Name Mat 206 April 15, 2019 →r-al °. .rag ,n a ry ok# -l:o
Answer #1

1) 5x^4 -10x^3 - 75x^2 = 0

dividing both sides by 5x^2

x^2 - 2x - 15 = 0

( x - 5) ( x+ 3) = 0

x = 5 , x = - 3

roots are x = 0 , - 3 , 5

2)

3x^4 - 21 x^3 + 10x^2 = 0

x^2 ( 3x^2 - 21x + 10 )= 0

x = 0

x = [ 21 +- sqrt ( 21^2 - 4810 * 3 )] / 2 *3

roots are x = 0 , ( 21 + sqrt 321)/6 , ( 21 - sqrt 321)/6